OpenStudy (anonymous):

The mean of 8 numbers is 56.25. What would the new mean be if the numbers 108 and 100 are added to the data set? 65.8 82.25 73.1 69.75

OpenStudy (anonymous):

@kevindragonlord @rosebud4612flvs

OpenStudy (anonymous):

it's the same thing.

OpenStudy (anonymous):

It's the same question as earlier -> the answer would be 65.8

OpenStudy (anonymous):

yeah its B

OpenStudy (perl):

we know that \[\frac {8*56.25}{8} = 56.25 \\ \frac {8*56.25 +100 + 108}{10} = \]